# Simply typed convertibility is TOWER-complete even for safe lambda-terms

### v2

@@ Abstract
1 Introduction
@@
1.1 Complexity of $\beta$-convertibility
@@
1.2 The case of the safe $\lambda$-calculus
@@
1.3 Some pedantic points
@@ 1.4 The $\eta$ rule
2 Simply typed $\beta$-convertibility is in Tower
2.1 A simple Tower normalization algorithm using the degree
@@
This definition on untyped $\lambda$-terms can be extended to Church-style simply typed $\lambda$-terms in the obvious way. It is immediate that $t\longrightarrow_{\beta}^{*}t^{\star}$. Furthermore:
Proposition 4 ([AL13, Appendix]).
Let $t$ be a simply typed $\lambda$-term whose redexes have degree at most $d\in\mathbb{N}$. Then:
@@
2.2 Finer bounds with the order
@@
3 Safe $\beta$-convertibility is Tower-hard
@@
3.1 Homogeneity and long-safety
@@
The second restriction is long-safety. In Church style, it is the same as the notion of safety in §1.2 except more subterms of a simply typed $\lambda$-term are considered to be in “unsafe position”: if $u$ is not a $\lambda$-abstraction, then it is in unsafe position in $t=C[\lambda x.\,u]$. Equivalently, in Curry style: {defi}[[BO09, Def. 1.21]] The typing rules of the long-safe $\lambda$-calculus are
@@
3.2 Star-free expressions
@@
3.3 From star-free expressions to safe $\lambda$-terms
@@
{defi}
[cf. [BB85, §4]] The type $\mathtt{Str}_{\Sigma}=\overbrace{(o\to o)\to\dots\to(o\to o)}^{|\Sigma|\ \text{times}}\to o\to o$ is the so-called Church encodings of strings. In the case of a unary alphabet, the type of Church numerals is $\mathtt{Nat}=\mathtt{Str}_{\{\mathrm{I}\}}=(o\to o)\to o\to o$.
@@ 3.5 Final remarks
3.5.1 On the complexity of star-free expression equivalence
@@
3.5.2 Safety in higher-order recursion schemes
@@
3.5.3 Implicit automata in typed $\lambda$-calculi and transducer theory
@@
Further works [Ngu21, PP24] in this “implicit automata in typed $\lambda$-calculi” research programme have focused on characterizations of string-to-string (or tree-to-tree) functions computed by transducers, i.e. automata with output. For reasons that are close to the aforementioned connection with HORS, as discussed in [Ngu21, §1.4.1], definability in the safe $\lambda$-calculus characterizes an important class of functions from 1980s transducer theory. The functions defined by the safe $\lambda$-terms $\mathtt{split}_{\Sigma}$ and $\mathtt{enum}_{\Sigma}$ (from Lemmas 13 and 14 respectively) therefore belong to this class. The “split” function is also a typical example of the more recently introduced polyregular functions, cf. [Boj22] – in fact, it is a slight variation on the “squaring with underlining” function of [Boj22, Example 3].

### v3

@@ Abstract.
Thanks to Anupam Das, Damiano Mazza and Noam Zeilberger for many instructive discussions on the topic of complexity of normalization for $\lambda$-calculi.
The author was supported by the LambdaComb project (ANR-21-CE48-0017) while working at École Polytechnique and by the LABEX MILYON (ANR-10-LABX-0070) of Université de Lyon, within the program “France 2030” operated by the French National Research Agency (ANR)
1. Introduction
@@
1.1. Complexity of $\beta$-convertibility
@@
1.2. The case of the safe $\lambda$-calculus
@@
1.3. Some pedantic points
@@ 1.4. The $\eta$ rule
2. Simply typed $\beta$-convertibility is in Tower
2.1. A simple Tower normalization algorithm using the degree
@@
This definition on untyped $\lambda$-terms can be extended to Church-style simply typed $\lambda$-terms in the obvious way. It is immediate that $t\longrightarrow_{\beta}^{*}t^{\star}$. Furthermore: {propC}[[AL13, Appendix]] Let $t$ be a simply typed $\lambda$-term whose redexes have degree at most $d\in\mathbb{N}$. Then:
@@
2.2. Finer bounds with the order
@@
3. Safe $\beta$-convertibility is Tower-hard
@@
3.1. Homogeneity and long-safety
@@
The second restriction is long-safety. In Church style, it is the same as the notion of safety in §1.2 except more subterms of a simply typed $\lambda$-term are considered to be in “unsafe position”: if $u$ is not a $\lambda$-abstraction, then it is in unsafe position in $t=C[\lambda x.\,u]$. Equivalently, in Curry style: {defiC}[[BO09, Def. 1.21]] The typing rules of the long-safe $\lambda$-calculus are
@@
3.2. Star-free expressions
@@
3.3. From star-free expressions to safe $\lambda$-terms
@@
{defiC}
[[BB85, §4]] The type $\mathtt{Str}_{\Sigma}=\overbrace{(o\to o)\to\dots\to(o\to o)}^{|\Sigma|\ \text{times}}\to o\to o$ is the type of so-called Church encodings of strings over the finite alphabet $\Sigma$. In the case of a unary alphabet, the type of Church numerals is $\mathtt{Nat}=\mathtt{Str}_{\{\mathrm{I}\}}=(o\to o)\to o\to o$.
@@ 3.5. Final remarks
3.5.1. On the complexity of star-free expression equivalence
@@
3.5.2. Safety in higher-order recursion schemes
@@
3.5.3. Implicit automata in typed $\lambda$-calculi and transducer theory
@@
Further works [Ngu21, PP24] in this “implicit automata in typed $\lambda$-calculi” research programme have focused on characterizations of string-to-string (or tree-to-tree) functions computed by transducers, i.e. automata with output. For reasons that are close to the aforementioned connection with HORS, as discussed in [Ngu21, §1.4.1], definability in the safe $\lambda$-calculus characterizes an important class of functions from 1980s transducer theory. The functions defined by the safe $\lambda$-terms $\mathtt{split}_{\Sigma}$ and $\mathtt{enum}_{\Sigma}$ (from Lemmas 12 and 13 respectively) therefore belong to this class. The “split” function is also a typical example of the more recently introduced polyregular functions, cf. [Boj22] – in fact, it is a slight variation on the “squaring with underlining” function of [Boj22, Example 3].